2. A 22.0 mL volume of 0.095 M NaOH is required to reach the phenolphthalein endpoint in

the titration of a 3.05 g sample of vinegar.

a. Calculate the number of moles of acetic acid in the vinegar sample.

b. Calculate the mass of acetic acid in the vinegar sample.

c. Calculate the percent by mass of acetic acid in the vinegar sample. Assume the density

of the vinegar is 1.0 g/mL.

3. In part B if the endpoint of the titration is overshot! Does this technique error result in an increase, a decrease, or have no effect on the reported percent acetic acid in the vinegar?

Explain.


1
Expert's answer
2020-12-01T09:05:11-0500

2.А

CH3COOH + NaOH --> CH3COONa + H2O

a) n(CH3COOH)= V(NaOH)*M(NaOH)=0.022*0.095=0.0021 mole;

b)m(CH3COOH)= n(CH3COOH)*Mr(CH3COOH)=0.0021*60=0.126 g;

c) % mass= 0.126/3.05 *100%= 4.13 %


3.The the percent acetic acid in the vinegar is directly proportional to the amount of NaOH added. It means that this technique error result in an increase on the reported percent acetic acid in the vinegar